### v1

@@ 6.1.2 Representation Learning
Representation learning involves training a parameterized model to learn the mapping from the raw input data to a feature vector, with the aim of uncovering more abstract and useful concepts. This process is designed to enhance the effectiveness of various downstream tasks by capturing the essential information embedded in the data [108]. In the context of DG, representation learning mainly focus on the concept of domain alignment for creating robust and generalized representations to unseen data. The goal of domain alignment is to minimize the difference among source domains for learning domain-invariant representations. It assumes that domain-invariant representation to the source domain should also be robust to unseen test domain. Recently, many methods have emerged to measure the distance between distributions and achieve domain alignment. These methods can be categorized into four main groups: adversarial learning, distance metric learning, contrastive learning, and regularization.
@@
Adversarial learning was initially proposed by Goodfellow et al. [62] to train a generative model, which can create realistic images. In essence, the training process consists of two neural networks competing against each other: a discriminator network and a generator network. The goal of the discriminator is to differentiate between images coming from the training data and the generated data. In contrast, the generator is tasked to generate fake data so realistic that can confuse the discriminator. In DG, adversarial learning is utilized to acquire source domain-invariant features that can be effectively used on new testing domains. In this setting, the learning goal is to confuse a domain discriminator, which is trained to detect the domains using features produced by the encoder.
@@
$\mathcal{L(C_{D})}=\frac{1}{B}\sum_{k=1}^{B}L(C_{D}(E(x_{k})),y_{k})$ (4)
where $x_{k}$ are samples drawn from source and target domain, $y_{k}$ are domain labels, with $y_{k}=1$ indicating $x_{k}$ comes from source domain and $y_{k}=0$ denoting $x_{k}$ comes from target domain. $L(.)$ is the cross-entropy loss. $B$ is the number of samples in the training batch.
In contrast, the category classifier is tasked to identify diseases. The category classifier is trained using the cross entropy loss:
$\mathcal{L(C_{S})}=\frac{1}{n}\sum_{i=1}^{n}L(C_{S}(E(x_{i})),y_{i})$ (5)
The end goal is to learn domain-invariant features across the source and target domain, that is to accurately predict labels without relying on any domain indicators. To this end, the feature extraction network is jointly trained to minimize the category classification loss (equation 5) and maximize the domain classification (equation 4) using stochastic gradient descent:
$\mathcal{L}_{total}=\mathcal{L(C_{S}})-\alpha\mathcal{L(C_{D})}$ (6)
where $\alpha$ is a hyperparameter to control the contribution of adversarial loss. In practice, given that the feature extraction network parameters are jointly updated by the backpropagation of the category classifier and domain discriminator, a self-defined gradient reversal layer is added to transmit negative gradient variations by the domain discriminator. Note that in the forward propagation, this layer acts as an identity transform.
By co-training the two classifiers, the model is supposed to learn domain-invariant features for both domains as well as strong classification performance for the source data, which increases the robustness of the learnt model when used for the target domain. This method has been evaluated on three benchmark neuroimaging datasets, namely, ADNI-1 [87], ADNI-2 [87], and AIBL [54] datasets, showing good results for brain dementia identification and disease progression prediction.
Similar to the previously mentioned approach, Wilm et al. [209] presented a domain adversarial RetinaNet approach [^8] https://github.com/DeepMicroscopy/MIDOG for object detection. They incorporated a gradient reversal layer and a domain classifier to the RetinaNet model. To learn domain invariant feature, the network is trained using the domain classification loss for all source domains, the bounding box regression loss and the instance classification loss. This method was applied for mitotic figure detection in the context of the MIDOG challenge. It achieved an F1 score of 0.7183 on the challenge’s test set.
To address the same problem, another work [123] proposed DAC R-CNN (section 6.1.1), an UDA technique based on adversarial training. Their network comprises a pretrained ResNet-50 as a backbone, three cascaded detection heads for high quality detector, and a PatchGAN discriminator. The role of the PatchGAN discriminator is to address the differences between different image domains like image color and style. It is trained to distinguish features between source images and target images. At the same time, the discriminator guides the training of the network in an adversarial manner. One advantage of PatchGAN is that it can be applied to images with arbitrary sizes. In comparison to [209], their model achieved a better F1 score of 0.7500 for mitosis detection on the preliminary test phase of the MIDOG challenge.
Janizek et al. [88] [^9] https://github.com/suinleelab/cxr_adv proposed an adversarial deconfounding approach for DG for pneumonia detection using chest radiographs. In particular, they proved improved out-of-hospital generalization performance of their model by making it invariant to the view position of chest radiographs (anterior-posterior vs. posterior-anterior). The classifier is trained to detect pneumonia from a chest radiograph. It is trained using an adversarial network, which is trained to determine the view from the output score of the classifier. Both networks are trained in an alternative way, first the adversary network is optimized then the classifier is trained to fool the adversary network and to detect pneumonia. This approach does not require data from the target domain, instead it is based on domain knowledge about the causal relationships involved in the data to identify nuisance variables. These variables might relate differently to the outcome in the test domain then in the source domain. To overcome this issue, adversarial technique has shown promising results to train a classifier that is invariant to the nuisance variable.
Chen et al. [34] also presented an adversarial framework, Cross-device and Cross-anatomy Adaptation Network (CCAN), for video analysis. They proposed to incorporate mutual information and multi-scale deep features to bridge the gap between high-end and low-end medical imaging devices and protocol. To overcome the problem of large anatomical variations, they used the concept of mutual information to align the distribution of multi-scale deep features in adversarial training. This was achieved by training two discriminators that focus on local and global representations. These discriminators are tasked with distinguishing whether pairs of features are related or unrelated based on their mutual information. In practice, the mutual information is computed using two binary classification losses, to detect whether two features are a positive or negative pair from the same image. The first loss, called local mutual information loss, aims to maximize mutual information between the local features (extracted from local convolutional feature maps) and the global semantic feature (the projected latent global feature representation). On the other hand, the second loss, the global mutual information loss, aims to maximize mutual information between the global semantic features, and the classifier predictions. Their approach showed promising results when evaluated for anatomy classification of ultrasound datasets. It showed significant improvement: the mean recognition accuracy increased by 20.8% and 10.0%, compared to a method without domain adaptation and a domain adaptation method based on adversarial learning (DANN), respectively.
@@
Feature distribution alignment An alternative to domain adversarial learning for achieving domain-invariant representations is to match the feature distributions. This is typically achieved using information theory based technique such as Kullback-Leibler (KL) divergence. By minimizing the KL divergence, all source domain representations are aligned with a Gaussian distribution. On the other hand, other strategies such as Minimax Entropy (MME) use the principle of entropy minimization to achieve domain alignment.
Meng et al. [132] [^10] https://github.com/qingjie99/MetFA proposed an UDA framework based on distance Metric guided Feature Alignment (MetFA) to learn domain-invariant latent representations. The main idea of MetFA is to align features between the source and target domain, in order to enhance the generalization of DL models. MetFA comprises three key steps: 1) a supervised classification on the source domain, 2) feature alignment, and 3) class distribution alignment.
First, a supervised model, which is composed of an encoder, a Gaussian embedding and a classifier, is trained on source labeled domain data using a standard classification loss (i.e., cross-entropy loss). Next, features of the source and target domains are aligned according to a predefined distribution within a shared latent space. That is, features from both domains are embedded into a common latent space, where they are guided to follow a predefined distribution (i.e., a Gaussian distribution). To this end, the KL divergence is employed to align the embeddings with a Gaussian distribution. In this shared latent space, the authors proposed to estimate class representations (prototypes) to constrain the features to be domain-invariant in each class.
In essence, this technique is inspired by MME [161], a semi-supervised DA technique that identifies domain-invariant prototypes and clusters target domain features around them. In MME, a single domain-invariant prototype (a representation point) is learned for each class. The distance between the class prototypes and neighboring unlabeled target samples is then minimized to extract discriminative features. The model architecture is composed of an encoding network and a classification layer, wherein the weight vectors are treated as the class-specific prototypes. The adaptation process in MME involves two steps. In the initial step, the prototypes are transitioned from the source domain to the target domain by maximizing the conditional entropy of unlabeled target data with respect to the classifier. This step is followed by minimizing the entropy with respect to the feature extractor to cluster features around the estimated prototype.
In addition, to enhance separation between clusters of different classes in both domains, they introduced cross-domain metric learning. It aims to minimize the distance between the latent features of target data ($z_{\mathcal{T}}^{f}$) and the latent features of labeled source data ($z_{\mathcal{S}}^{f}$) when they belong to the same class, while maximizing the distance when they are from different classes. The metric loss is defined as follows:
$\begin{split}\mathcal{L}&=\frac{1}{n_{\mathcal{T}}}\sum_{i=1}^{c}\sum_{j=1}^{c_{\mathcal{T}i}}log(1+\sum_{k\neq i}^{k\in[1,c]}e^{d_{j}^{i}-d_{j}^{k}})\end{split}$ (7)
where $c$ is the number of classes. $n_{\mathcal{T}}$ and $c_{\mathcal{T}i}$ are the number of all latent features of target samples and the number of latent features of target samples from class $i$.
Using a hard mining strategy [32], $d_{j}^{i}$ is the distance between $z_{\mathcal{T}j}^{fi}$ and a same class $z_{\mathcal{S}t}^{fi}$:
$d_{j}^{i}=\max_{t}d(z_{\mathcal{T}j}^{fi},z_{\mathcal{S}t}^{fi}),t\in[1,c_{\mathcal{S}i}]$ (8)
where $c_{\mathcal{S}i}$ denotes the number of $z_{\mathcal{S}}^{f}$ from class $i$ and $d(.,.)$ the squared Euclidean distance.
$d_{j}^{k}$ is the distance between $z_{\mathcal{T}j}^{fi}$ and a different class $z_{\mathcal{S}t}^{fk}$:
$d_{j}^{k}=\min_{t}d(z_{\mathcal{T}j}^{fi},z_{\mathcal{S}t}^{fk}),t\in[1,c_{\mathcal{S}k}]$ (9)
where $c_{\mathcal{S}k}$ denotes the number of $z_{\mathcal{S}}^{f}$ from class $k$.
Beyond aligning individual features, MetFA also aligns the class distributions between the source and target domains. To this end, similar to Dou et al. [50], a soft label distributions with a "softened" softmax at temperature $\tau$ is computed for the target and the source domain. Specifically, for the source domain, the soft label distribution ($s_{\mathcal{S}i}$) is computed as follows: $s_{\mathcal{S}i}=softmax(\frac{1}{\tau_{1}}\frac{1}{c_{\mathcal{S}i}}\sum_{y=i}z_{\mathcal{S}}^{c}\ )|_{(x,y)\sim\{\mathcal{S}\}}$, where $z_{\mathcal{S}}^{c}$ are the pre-softmax activations from the classifier for the source domain samples and $\tau_{1}$ is a temperature parameter. Similarly, $s_{\mathcal{T}i}$ is the $i^{th}$ class distributions in target domain such as: $s_{\mathcal{T}i}=softmax(\frac{1}{\tau_{1}}\frac{1}{c_{\mathcal{T}i}}\sum_{y=i}z_{\mathcal{T}}^{c}\ )|_{(x,y)\sim\{\mathcal{T}\}}$, $z_{\mathcal{T}}^{c}$ are the pre-softmax activations from the classifier for the target domain samples and $c_{\mathcal{T}i}$ is the number of target domain samples predicted for class $i$. The symmetrized KL divergence is then used to define the class distribution alignment loss between $s_{\mathcal{S}i}$ and $s_{\mathcal{T}i}$.
[…]
@@
Contrastive learning Domain alignment based on contrastive learning has emerged as an effective strategy. These techniques often employ contrastive learning, a machine learning paradigm which can be viewed as learning by comparing. In contrastive learning, a representation is learned by comparing among the input samples. The comparison can be conducted between pairs of similar inputs (positive pairs) and pairs of dissimilar inputs (negative pairs). The method involves computing the distance between feature vectors of image pairs and deriving the loss according to this distance. An image pair is deemed sufficiently dissimilar if the computed distance exceeds a predetermined margin. Typically, most studies considered images coming from the same class as similar pairs.
@@
Gurpinar et al. [72] proposed a DA approach based on contrastive learning with cosine distance to align data distributions between source and target domain. Their approach consists of using the contrastive learning framework to align the embedding features, with the aim to decrease the discrepancy between the source and target domains. While the definition of positive and negative pairs is considered to be explicit in multi-class classification problem, in multi-label problem it becomes more challenging to form these pairs. To this end, Gurpinar et al. [72] used a Siamese network to learn embeddings such that samples from the same class are gathered closer and samples from different classes are pushed away. In addition, they added a new smoothing parameter, $\beta$, to the loss to make it proportional to the similarity regarding present labels.
@@
where $d$ denotes the cosine distance between feature vectors extracted from the pair of images, $m$ indicates the margin and $Y$ represents the pair labels ($Y=1$ for dissimilar pairs and $Y=0$ for similar pairs). $\beta$ values range between 0 and 1.
They applied this method for building a facial action unit detection model for children with hearing impairments, showing that integrating $\beta$ leads to improved results between and 3% to 14%.
Le et al. [107] argue that data augmentation allows the extension of the training data but does not provide a theoretical guarantee for DG. On the other hand, invariant feature representation learning offers theoretical guarantee but the number of targeted domains that can be addressed is limited by the original training domains. To boost the performances, they proposed to combine both approaches. In particular, they applied data augmentation including random cropping, random horizontal flip, random color jitter, random grayscale, random Gaussian blur, normalization, and random erasing. To enforce invariant features, they proposed to minimize the distance between original and augmented domains using a supervised contrastive learning loss. Specifically, they adopted the contrastive loss in the form of normalized temperature-scaled cross-entropy loss (NT-Xent). The essence of this method is to minimize the distance between original and augmented domains. Positive pairs consist of pair of samples from the same class while negative pairs are from different classes.
The feature extractor was trained to optimize the cross-entropy loss and the contrastive loss while the classifier was optimized to minimize the cross-entropy loss. They used a different learning rate to update the feature extractor and the classifier. A disadvantage of this method is it assumes that the label distribution to be roughly equal across the domains, implying the need for balanced datasets. It was assessed on the medical benchmark dataset of chest X-ray dataset and on DG benchmark PACS.
@@
Consistency regularization Consistency regularization methods in representation learning mainly focus on adding a loss to the learning objective for making the model robust against variations in input data that are irrelevant to the classification task such as changes in texture, color, etc.
With the aim to develop generalized representations, Zhang et al. [231] proposed a semi-supervised DG method, called Stable and Orthogonal Feature Regularization (SOFR). The main idea is to make the learned representation have two characteristics: stability and orthogonality. The stability enforces the features to be domain-invariant using a data augmentation technique, domain randomization (see Section 6.1.1), while the orthogonality avoids overfitting by enforcing minimal redundancy.
@@
To enforce the model to learn domain-invariant features that are orthogonal to each other, two regularization factors are integrated to the learning objective. For enforcing feature stability, a regularization based on an instance-level paradigm is used. In particular, it consisted of computing the sum of the channel-wise cosine distance between the original feature and its augmented version (its domain-randomized counterpart). The orthogonality of features is measured by cosine similarity between different channels. This method was applied for Chest X-ray diagnosis, using MIMIC [90] as the labeled source and NIH [207] and CXP [85] as unlabeled source domain. It was tested on the PadChest [28] dataset, showing promising results with a mean AUC of 0.8443.
Raipuria et al. [151] proposed a regularization loss to achieve better generalization performances in computational histopathology. They introduce a consistency regularization loss to constrain their model to achieve high invariance to stain color changes on unseen test data. Using this technique, the model is enforced to produce consistent predictions between original samples and their stain modified versions. This is done by minimizing the KL divergence loss:
$\mathcal{L}_{consistency}=D_{KL}(C(F(E(x)))\ ||\ C(F(E(\hat{x}))))$ (12)
where $C$ denotes the classifier, $F$ is the projector, $E$ is the encoder, $\hat{x}$ is the stained altered image, $D_{KL}$ is the symmetrized KL divergence.
@@
In addition to the consistency regularization loss, stain regeneration loss is employed to train a decoder to regenerate the original stain color using feature representation of the stain modified images ($\hat{x}$). To this end, they employed the mean squared error loss between the raw image and the regenerated image. This loss is considered as an auxiliary task which helps the model to improve generalization on images with stain variation by learning shared features for classification and regeneration task. This method was evaluated on two publicly available datasets TUPAC Mitosis Detection and Camelyon17 Tumor Metastasis Detection. It showed that stain-invariant features results in improved performance on unseen images coming from different centers.
For the purpose of DG in real-time surgical tool detection, Reiter [157] proposed to improve a transformer-based architecture, DEtection TRansformer (DETR) [29], by adding a latent feature space using variational encoding. This step is designed to capture the common intra-domain information by fitting the latent features to a prior distribution, a multivariate Gaussian. In addition, they employed a rank regularization to enhance the model’s ability to generalize across different domains. This approach is based on the assumption that there are linear dependencies between the latent features of different domains. This means that the latent features from one domain can be linearly related to those of another domain. In particular, it is assumed that a matrix formed from the latent features for a given label contains an influential eigenvalue related to that label. This allows the regularization of intra-class variability by constraining the rank of the latent features to the number of classes. This is done by minimizing the corresponding diagonal value in the matrix returned by singular value decomposition. The results of this method has shown improved DG performance for end-to-end detection in surgical tool detection, when assessed on three public datasets showing different instrument types and procedures.
Similarly, Li et al. [112] proposed a Linear-Dependency Domain Generalization (LDDG) [^11] https://github.com/wyf0912/LDDG method to improve the generalizability of DL model for medical image classification. Similar to the aforementioned method, the authors used a linear-dependency modeling and imposed a rank regularization based on the latent feature space. Moreover, they used variational encoding for distribution alignment. To this end, they adopt KL divergence for matching the latent features from multiple source domains to a pre-defined prior distribution, which in this paper was defined as Gaussian distribution. They evaluated this approach on two challenging medical imaging classification tasks: skin lesion classification and Spinal cord gray matter segmentation.
@@
Viviano et al. [200] investigated whether poor generalization in DL models can be visually diagnosed using attribution (or saliency) maps [^12] https://github.com/josephdviviano/saliency-red-herring. The motivation behind this is that saliency maps could demonstrate if the DL model is behaving unexpectedly. For instance, instead of learning robust representation, i.e., features that denote a class, the DL model can learn spuriously-correlated image features present only in the training distribution. This behavior is also known as shortcut learning or incorrect feature attribution, where the model overfits to a set of training data. The authors hypothesized that models showing good generalization properties have attribution maps which only use the class-discriminative features to make predictions. To this end, they proposed a regularization strategy where the DL model is trained to ignore confounders (such as acquisition site or other image acquisition parameters) using attribution priors, i.e., to produce predictions using the correct anatomy (as a medical expert would). To accomplish this, they used medical datasets with auxiliary labels which consisted of masks drawn by an expert, indicating a region of the image containing relevant information to make the predictions. To make the model invariant to features arising from outside of the attribution priors, they studied two feature-aware-methods: 1) the Activation Difference (ActDIff) Approach, which regularizes the model and penalizes the L2-normed distance between the masked and unmasked input’s latent representations, 2) adversarial approach which employs a discriminator to identify whether latent representations come from a masked or unmasked input. In parallel, the encoder is trained to fool the discriminator and produce discriminative representations. In addition, they explored the saliency penalties, where GradMask [181] and Right for the Right Reasons (RRR)[159] are applied on the input space. The idea is to penalize the model for producing saliency gradients outside of regions of interest denoting the class-discriminative feature. The results showed that although the DL models were explicitly regularized to not use features constructed from outside of the mask, the DL network still attribute features outside of the mask at test time. Therefore, the authors concluded that the proposed methods do not guarantee to negate any confounding variables that exist within the mask.

### v3

@@ 5.1.2 Representation Learning
Representation learning involves training a parameterized model to learn the mapping from the raw input data to a feature vector, with the aim of uncovering more abstract and useful concepts. This process is designed to enhance the effectiveness of various downstream tasks by capturing the essential information embedded in the data [121]. In the context of DG, representation learning mainly focus on the concept of domain alignment for creating robust and generalized representations to unseen data. The goal of domain alignment is to minimize the difference among source domains for learning domain-invariant representations. It assumes that domain-invariant representation to the source domain should also be robust to unseen test domain. Recently, many methods have emerged to measure the distance between distributions and achieve domain alignment. These methods can be categorized into four main groups: adversarial learning, feature distribution alignment, contrastive learning, and consistency regularization.
@@
In DG, adversarial learning is utilized to acquire source domain-invariant features that can be effectively used on new testing domains. In general, this is achieved by training an encoder ($E$) with an adversary discriminator (i.e., a domain discriminator, $C_{D}$) and a category classifier ($C_{S}$), as illustrated in Figure 5. The domain discriminator is tasked to distinguish the domains of the input features by minimizing the cross-entropy loss (${\mathcal{L}_{d}}$). The category classifier is employed for the main task of classification (${\mathcal{L}_{c}}$). The end goal is to learn domain-invariant features across the source domains, that is to accurately predict disease labels without relying on any domain indicators. To this end, the feature extraction network is trained to confuse a domain discriminator and to accurately classify diseases. It is jointly trained to maximize the domain classification loss and to minimize the category classification loss:
$\mathcal{L}_{total}=\mathcal{L}_{c}-\alpha\mathcal{L}_{d}$ (4)
where $\alpha$ is a hyperparameter to control the contribution of adversarial loss. In practice, given that the feature extraction network parameters are jointly updated by the backpropagation of the category classifier and the domain discriminator, a self-defined gradient reversal layer is added to transmit negative gradient variations from the domain discriminator. Note that in the forward propagation, this layer acts as an identity transform.
Figure 5: Adversarial learning.
For instance, Wilm et al. [45] incorporated a gradient reversal layer and a domain classifier to their RetinaNet model [^7] https://github.com/DeepMicroscopy/MIDOG employed for object detection. To learn domain invariant feature, the network is trained using the domain classification loss for all source domains, the bounding box regression loss and the instance classification loss. For the task of mitotic figure detection, this method achieved an F1 score of 0.7183 on the MIDOG challenge’s test set.
To address the same problem, another work [38] proposed an UDA technique based on adversarial training. It comprised a pretrained ResNet-50 as a backbone, three cascaded detection heads for high quality detector, and a PatchGAN discriminator. The PatchGAN discriminator is trained to distinguish features between source images and target images. At the same time, it guides the training of the network in an adversarial manner. One advantage of PatchGAN is that it can be applied to images with arbitrary sizes. In comparison to [45], their model achieved a better F1 score of 0.7500 for mitosis detection on the preliminary test phase of the MIDOG challenge.
Guan et al. [46] also presented an UDA approach based on adversarial learning, named Attention-driven Deep Domain Adaptation. It is composed of: 1) a feature encoding module, 2) an attention discovery module that discovers disease-related regions and 3) a domain transfer module with adversarial learning comprising two classifiers: a domain discriminator and a category clasifier. By co-training the two classifiers, the model is supposed to learn domain-invariant features for both domains (source and target domains) as well as strong classification performance for the source data, which increases the robustness of the learnt model when used for the target domain. This method showed good results for brain dementia identification and disease progression prediction when evaluated on three benchmark neuroimaging datasets ADNI-1 [122], ADNI-2 [122], and AIBL [123] datasets.
In their UDA adversarial framework, Cross-device and Cross-anatomy Adaptation Network, Chen et al. [47] aimed to enhance anatomy classification in ultrasound video. Their main idea was to align the distribution of multi-scale deep features in adversarial training. This alignment involved training two discriminators, a local and a global discriminator, that assess whether pairs of features are a positive or negative pair from the same image based on their mutual information. The local discriminator enhances the correlation between local convolutional features and a unified global semantic feature, while the global discriminator aligns the global semantic feature with the classifier predictions. Their approach showed promising results in ultrasound anatomy classification, with mean recognition accuracy increasing by 20.8% and 10.0%, compared to a method without domain adaptation and an adversarial learning-based domain adaptation method, respectively.
Janizek et al. [48] [^8] https://github.com/suinleelab/cxr_adv proposed an adversarial deconfounding approach to improve pneumonia detection in chest X-rays. Their aim was to make their pneumonia detection model invariant to the view position of chest radiographs (anterior-posterior vs. posterior-anterior). This was achieved by jointly training a classifier with an adversarial network that tries to determine the view from the classifier output score. In contrast to previously mentioned methods, this approach does not require data from the target domain, instead it is based on domain knowledge about the causal relationships involved in the data to identify nuisance variables. These variables might relate differently to the outcome in the test domain then in the source domain. To overcome this issue, adversarial technique has shown promising results to train a classifier that is invariant to the nuisance variable.
@@
Feature distribution alignment An alternative to domain adversarial learning for achieving domain-invariant representations is to match the feature distributions. This is typically achieved using information theory based technique such as Kullback-Leibler (KL) divergence. By minimizing the KL divergence, all source domain representations are aligned with a Gaussian distribution. On the other hand, other strategies such as Minimax Entropy (MME) [124] use the principle of entropy minimization to achieve domain alignment.
In their UDA framework, MetFA [^9] https://github.com/qingjie99/MetFA, Meng et al. [49] proposed to learn a shared latent representation space between the source and target domains using a Gaussian embedding modeled by a standard Gaussian distribution. This distribution matching is achieved through the KL divergence. Inspired by MME [124], class representations (prototypes) are estimated in this shared latent space. These prototypes, which correspond to the weights of the last dense layer in the classifier, are initially transitioned from the source domain to the target domain by maximizing the conditional entropy of unlabeled target data. In the second step, features are clustered around these prototypes by minimizing the entropy with respect to the feature extractor. Furthermore, in order to maximize the margin between different classes across domains, a cross-domain metric learning was proposed. It aims to minimize the distance between the latent features of the target data (query samples) and the latent features of the labeled source data (support samples) when they belong to the same class, while maximizing the distance when they are from different classes.
Additionally, MetFA aligns the class distributions between the source and target domains. Following Dou et al. [125], soft label distributions are computed for both domains using a “softened” softmax at temperature $\tau$. The class distribution alignment loss is then assessed using the symmetrized KL divergence between these soft label distributions. This method was evaluated for cross-device anatomical classification of fetal ultrasound view planes. It achieved an F1 score of 0.5776 and of 0.7713 on the target data coming from the GE Voluson E8 device and the Philips EPIQ V7 G device, respectively.
@@
Contrastive learning Domain alignment based on contrastive learning has emerged as an effective strategy. These techniques often employ contrastive learning, a machine learning paradigm which can be viewed as learning by comparing. In Contrastive Learning (CL), a representation is learned by comparing among the input samples. The comparison can be conducted between pairs of similar inputs (positive pairs) and pairs of dissimilar inputs (negative pairs). The method involves computing the distance between feature vectors of image pairs and deriving the loss according to this distance. An image pair is deemed sufficiently dissimilar if the computed distance exceeds a predetermined margin. Typically, most studies considered images coming from the same class as similar pairs.
@@
Gurpinar et al. [50] proposed a DA approach based on contrastive learning with cosine distance. A Siamese network is trained using a CL loss to learn embeddings such that samples from the same class are gathered closer and samples from different classes are pushed away. To further adapt it to a multi-label classification problem, a new smoothing parameter, $\beta$ (ranging between 0 and 1), is added to the loss to make it proportional to the similarity regarding present labels.
@@
where $d$ denotes the cosine distance between feature vectors extracted from the pair of images, $m$ indicates the margin and $Y$ represents the pair labels ($Y=1$ for dissimilar pairs and $Y=0$ for similar pairs). This method was applied for facial action unit detection for children with hearing impairments. Integrating $\beta$ led to improved recognition performance with a weighted F1 score ranging between 0.76 and 0.85 on the target HIC dataset.
Le et al. [51] combined data augmentation approaches with domain alignment based on CL. The essence of this method is to minimize the distance between original and augmented domains. Positive pairs consist of pair of samples from the same class while negative pairs are from different classes. Augmented domains were obtained by applying techniques such as random cropping, random horizontal flip, random color jitter, and random grayscale. To enforce invariant features, the distance between original and augmented domains was minimized using a supervised contrastive learning loss in the form of normalized temperature-scaled cross-entropy loss. This method was assessed on PACS and on a medical benchmark dataset of chest X-ray, consisting of data from CheXpert, ChestX-ray14 and PadChest. A disadvantage of this method is it assumes that the label distribution to be roughly equal across the domains, implying the need for balanced datasets.
@@
Consistency regularization Consistency regularization methods mainly add a loss term to the learning objective to make the model robust against variations in input data that are irrelevant to the classification task such as changes in texture, color, etc.
@@
Zhang et al. [20] proposed a semi-DG method, which constrains the learned representation to have two characteristics: stability and orthogonality. Their regularization objective was applied to features from (labeled or unlabeled) pairs of original and domain-randomized augmented images. To enforce feature stability, the sum of the channel-wise cosine distance between the original feature and its augmented version (its domain-randomized counterpart) was computed. The orthogonality of features was assessed using the cosine similarity between different channels. This method was applied for chest X-ray diagnosis, using MIMIC [126] as the labeled source and NIH [127] and CXP [128] as unlabeled source domain. It was tested on the PadChest [129] dataset, showing promising results with a mean AUC of 0.8443 for detecting pathologies.
@@
In the context of computational histopathology, Raipuria et al. [52] introduced a consistency regularization loss to ensure their model remains highly invariant to stain color changes on unseen test data. Their model was enforced to produce consistent predictions for both original samples and their stain modified versions using the KL divergence loss. In addition, an auxiliary task of stain regeneration was applied to enhance the model’s generalization capabilities. This involves training a decoder to regenerate the original stain color using feature representation of the stain modified images ($\hat{x}$). Thereby, a shared representation is learned for the primary task of classification and the auxiliary task of stain regeneration. This method was evaluated on two publicly available datasets TUPAC-16 and Camelyon17. It showed that stain-invariant features results in improved performance on unseen images coming from different centers.
To improve DG performances, Li et al. [53] presented a rank-regularized latent feature space [^10] https://github.com/wyf0912/LDDG. Based on the assumption that there are linear dependencies between the latent features of different domains, the latent feature space was regularized by modeling intra-class variation using rank constraint: the rank of the latent feature matrix was constrained to the number of classes. At the same time, the distribution of latent features was aligned to a common Gaussian distribution. This approach was evaluated for skin lesion classification using seven public skin lesion datasets. Using ResNet18 model, it showed better cross-domain generalization performances when compared to state-of-the-art baselines. Inspired by these results, Reiter [54] incorporated this approach in their Detection Transformer (DETR)  [^11] https://github.com/facebookresearch/detr model [130] for the purpose of DG in real-time surgical tool detection. Despite reporting improved generalization performances, the limited size of the datasets was a major limitation.
@@
Viviano et al. [55] explored different regularization strategies [^12] https://github.com/josephdviviano/saliency-red-herring where the DL model is trained to ignore confounders (such as acquisition site) using attribution (saliency) priors, i.e., expert-drawn masks highlighting relevant regions for predictions. These methods consisted of: 1) an activation difference approach, which regularizes the model and penalizes the L2-normed distance between the masked and unmasked input’s latent representations, 2) an adversarial approach which employs a discriminator to identify whether latent representations come from a masked or unmasked input, and 3) two saliency penalties methods (GradMask [131] and Right for the Right Reasons (RRR)[132]) which penalize the model for producing saliency gradients outside of regions of interest. Despite improved generalization performance in the presence of covariate shift, the results showed that the DL network still attribute features outside of the mask at test time. Indeed, the proposed methods do not guarantee to negate any confounding variables that exist within the mask.